# Debiasing surgeon: fantastic weights and how to find them

### v1

@@ B.1 Is FFW helping the network focus on the right features ?
B.2 Pruning distribution across the networks.
(a) Biased MNIST on a small CNN.
(c) Corrupted CIFAR10 on ResNet18.
@@
Appendix C More ablations
C.1 Variations on the Mutual Information Minimization
Figure 5: Results for FFW applied on Biased-MNIST ($\rho=0.99$) with different $\gamma$ on the validation set.
@@ C.1 Variations on the Mutual Information Minimization
C.2 Comparison of Pruning Strategies
Table 7: Results for Biased-MNIST ($\rho=0.99$ with different pruning strategies.
@@ C.2 Comparison of Pruning Strategies
C.3 Effect of the extraction point
@@
Appendix D Details for Section 3
In this section, we will provide all the derivations for Sec. 3 in the main paper. more specifically, Sec. D.1 will propose the derivation for (3), Sec. D.2 will discuss the joint probability that will be employed to derive (4) in Sec. D.3, and (6) in Sec. D.4.
D.1 Derivation for (3)
@@
finding back (3).
D.2 Joint probability between $\hat{B}$, $\hat{Y}$, $Y$
A clear dependency between $\rho$ and (3), as already showcased in Sec. D.1, exists. This measure is applied to ground-truth labels, investigating the common information between them (and for instance, the information that is possible to disentangle). Nonetheless, in the more general case, the trained model (whose output is modelizable as the random variable $Y$) is not a perfect learner, having $H(Y|\hat{Y})\neq 0$. The model, in this case, does not correctly classify the target for two reasons.
@@ D.3 Derivation for (4)
$=$ $\frac{\rho+\phi(1-\rho)}{\log_{2}(N)}\log_{2}\left[\rho+\phi(1-\rho)\right]+\frac{(1-\phi)(1-\rho)}{\log_{2}(N)}\log_{2}\left[\frac{(1-\phi)(1-\rho)}{N-1}\right]+1,$ (15)
finding back (4).
D.4 Derivation for (6)
Similarly to the approach taken in Sec. D.3, from the joint probability as in (12), we marginalize, but this time on $\hat{B}$, by summing all the $N_{B}$ (in our simplified case, $N$) biases per given target class and prediction of the model. Under the assumption that $\varepsilon=\varepsilon_{\text{bia}}$, we have:

### v2

@@ 0.A.1 Is FFW helping the network focus on the right features ?
0.A.2 Pruning distribution across the networks.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
0.A.3 Variations on the Mutual Information Minimization
Figure 5: Results for FFW applied on Biased-MNIST ($\rho=0.99$) with different $\gamma$ on the validation set
@@ 0.A.3 Variations on the Mutual Information Minimization
0.A.4 Comparison of Pruning Strategies
Table 7: Results for Biased-MNIST ($\rho=0.99$) with different pruning strategies.
@@ 0.A.4 Comparison of Pruning Strategies
0.A.5 Effect of the extraction point
@@
0.A.6 Training times
We provide the average training times of our method (after training the vanilla model) in Tab. 8.
Table 8: Training time for our main experiments performed on a NVIDIA A100-PCIE-40GB GPU.
Dataset Model used Training time Unbiased training set
Biased-MNIST [4] SimpleConvNet [4] 3h20 36000
CelebA [38] ResNet18 1h15 437
Corrupted CIFAR10 [25] ResNet18 0h52 6000
Multi-Color MNIST [37] 3-layers MLP 0h22 6000
This duration varies with the sizes of the unbiased training set, of the input images as well as the model type.
Appendix 0.B Details for Section 3
In this section, we will provide all the derivations for Sec. 3 in the main paper. more specifically, Sec. 0.B.1 will propose the derivation for (3), Sec. 0.B.2 will discuss the joint probability that will be employed to derive (4) in Sec. 0.B.3 and (6) in Sec. 0.B.4.
0.B.1 Derivation for (3)
@@
0.B.2 Joint probability between $\hat{B}$, $\hat{Y}$, $Y$
A clear dependency between $\rho$ and (3), as already showcased in Sec. 0.B.1, exists. This measure is applied to ground-truth labels, investigating the common information between them (and for instance, the information that is possible to disentangle). Nonetheless, in the more general case, the trained model (whose output is modelizable as the random variable $Y$) is not a perfect learner, having $H(Y|\hat{Y})\neq 0$. The model, in this case, does not correctly classify the target for two reasons.
@@ 0.B.3 Derivation for (4)
$=$ $\frac{\rho+\phi(1-\rho)}{\log_{2}(N)}\log_{2}\left[\rho+\phi(1-\rho)\right]+\frac{(1-\phi)(1-\rho)}{\log_{2}(N)}\log_{2}\left[\frac{(1-\phi)(1-\rho)}{N-1}\right]+1.$ (15)
We apologize for the typos in (4), which will be corrected as in (15) in the final version of the paper.
0.B.4 Derivation for (6)
Similarly to the approach taken in Sec. 0.B.3, from the joint probability as in (12), we marginalize, but this time on $\hat{B}$, by summing all the $N_{B}$ (in our simplified case, $N$) biases per given target class and prediction of the model. Under the assumption that $\varepsilon=\varepsilon_{\text{bia}}$, we have: